OpenStudy (wilder.monday):

PLEASE HELP 1. 2^x-1 = 5^2x+6

OpenStudy (wilder.monday):

@zepdrix

OpenStudy (anonymous):

is it \[2^{x*1}=5^{2x+6}\]?

OpenStudy (wilder.monday):

yes :)

OpenStudy (wilder.monday):

x-1 actually

OpenStudy (anonymous):

ok be prepared to do some algebra

OpenStudy (wilder.monday):

oh goodie

OpenStudy (anonymous):

you gotta start with \[(x-1)\ln(2)=(2x+6)\ln(5)\] as a first step

OpenStudy (anonymous):

don't be confused by this \(\ln(2)\) and \(\ln(5)\) are just some numbers

OpenStudy (anonymous):

constants in other words, we are solving for \(x\) so the first thing you do is to distribute on both sides of the equal sign

OpenStudy (wilder.monday):

i think i get it

OpenStudy (anonymous):

we can take an additional step at the beginning, but it is unnecessary the math teachers say "take the log" and write \[\ln(2^{x-1})=\ln(5^{2x+6})\] but then the next step is to rewrite that (the whole point is to bring the \(x\) out of the sky) and write \[(x-1)\ln(2)=(2x+6)\ln(5)\]

OpenStudy (anonymous):

ready for the algebra?

OpenStudy (wilder.monday):

oh i get it now. go on please :)

OpenStudy (anonymous):

distribute first, remembering ( i repeat) that \(\ln(2)\) and \(\ln(5)\) are just numbers so first second step is to write \[\ln(2)x-\ln(2)=2\ln(5)x+6\ln(5)\] let me know when you digest that

OpenStudy (wilder.monday):

on the left side is it - ln(2) because in x - 1 the one has a - in front ?

OpenStudy (anonymous):

it is the distributive property in action \((b+c)a=ab+ac\) here \(b=x.c=-1,a=\ln(2)\) so \[(x-1)\ln(2)=\ln(2)\times x-1\times \ln(2)\]

OpenStudy (anonymous):

more succinctly written as \[\ln(2)x-\ln(2)\] same story on the other side

OpenStudy (wilder.monday):

okay so what would the next step be?

OpenStudy (anonymous):

you are solving for \(x\) so you have to get everything with an \(x\) on one side of the equal sign, everything else on the other

OpenStudy (anonymous):

it doesn't matter which side you pick but lets put the \(x\) stuff on the left ok?

OpenStudy (anonymous):

again don't be confused the the logs, it is like solving \[Ax-A=2Bx+6B\] for \(x\)

OpenStudy (wilder.monday):

okay so i do the opposite operation right which would be subtracting 2ln(5)x ...?? or am i really off

OpenStudy (anonymous):

no you are not off (not that i know of anyway) you are exactly right

OpenStudy (wilder.monday):

okay phew

OpenStudy (anonymous):

\[\ln(2)x-\ln(2)=2\ln(5)x+6\ln(5)\\ \ln(2)x-2\ln(5)x-\ln(2)=6\ln(5)\] after subtracting

OpenStudy (wilder.monday):

yes i got that right!

OpenStudy (anonymous):

now we got to get the term without the \(x\) on the other side

OpenStudy (wilder.monday):

yay!

OpenStudy (anonymous):

yay indeed!

OpenStudy (anonymous):

so since you are on a roll, what do you do now to get the term without an x on the right?

OpenStudy (wilder.monday):

so + in(2) on both sides

OpenStudy (anonymous):

bingo

OpenStudy (wilder.monday):

and from there...

OpenStudy (anonymous):

\[\ln(2)x-2\ln(5)x=6\ln(5)+\ln(2)\] is where we should be

OpenStudy (wilder.monday):

yes :)

OpenStudy (anonymous):

ok now here is the only part that requires some sophistication before we do it, could you solve \[Ax+Bx=C\] for \(x\) if you had to ? "no" is a fine answer, just asking

OpenStudy (wilder.monday):

for which side?

OpenStudy (anonymous):

solve for \(x\)

OpenStudy (anonymous):

if not, that is fine, i will show you

OpenStudy (wilder.monday):

yes please

OpenStudy (anonymous):

ok first imagine we had to solve \[3x+7x=15\] we would add first and get \[10x=15\] then divide by 10

OpenStudy (wilder.monday):

yes i understand that

OpenStudy (anonymous):

if we have \[Ax+Bx=C\] we can't really add (or combine like terms) but we would do the same thing \[(A+B)x=C\] so \[x=\frac{C}{A+B}\]

OpenStudy (anonymous):

in other words, factor out the \(x\) so we know what to divide by here we have \[\ln(2)x-2\ln(5)x=6\ln(5)+\ln(2)\] so we have to factor out the \(x\) on the left and get \[(\ln(2)-2\ln(5))x=6\ln(5)+\ln(2)\]

OpenStudy (wilder.monday):

ooooh yes i see

OpenStudy (anonymous):

let me know if that last step is clear, it is the only part of this that is confusing

OpenStudy (anonymous):

i going to take that as a "i get it" final step is to divide to finish with \[x=\frac{6\ln(5)+\ln(2)}{\ln(2)-2\ln(5)}\]

OpenStudy (wilder.monday):

yes i understand

OpenStudy (anonymous):

told you it was a bunch of algebra the only not algebra part was taking the log at the beginning to get the variable on the ground floor

OpenStudy (anonymous):

want to check it?

OpenStudy (wilder.monday):

im doing the calculations now :)

OpenStudy (anonymous):

here is an easy check http://www.wolframalpha.com/input/?i=+2^%28x-1%29+%3D+5^%282x%2B6%29 you can see that it is right

OpenStudy (anonymous):

now i have a question

OpenStudy (anonymous):

what the monkey is a "wilder monday"?

OpenStudy (wilder.monday):

haha i get that a lot

OpenStudy (wilder.monday):

its actually my name. Wilder-Monday Turner is my full name

OpenStudy (wilder.monday):

i have no idea why my mother did that but people just call me Monday

OpenStudy (anonymous):

that is a pretty cool name i know someone named "tuesday wolf" which i also like

OpenStudy (wilder.monday):

thats a nice name!

OpenStudy (anonymous):

and not that you asked, but "satellite73" is my car http://www.dvap.com/salvage/1973-plymouth-satellite-73pl5238d/ mine is nicer but similar

OpenStudy (wilder.monday):

i feel bad for not asking ahaha. thats a sick ride you got there

OpenStudy (anonymous):

lol don't feel bad you got more or are you done?

OpenStudy (wilder.monday):

more questions? yes i have more if you would like to help me :)

OpenStudy (anonymous):

sure why not? it is nice to help someone who actually participates too

OpenStudy (wilder.monday):

i like being involved because it helps me understand better instead of having someone do it all for me

OpenStudy (wilder.monday):

ill make a new question or do you want to keep going on this one?

OpenStudy (anonymous):

you can go ahead an post here if you like

OpenStudy (wilder.monday):

okay :) 1. 3 + e^3x+2 = 23

OpenStudy (anonymous):

it is \[1.3+e^{3x+2}=23\]?

OpenStudy (wilder.monday):

No no its just 3

OpenStudy (wilder.monday):

sorry bout that

OpenStudy (anonymous):

ooh \[3+e^{3x+2}=23\]

OpenStudy (wilder.monday):

yes yes

OpenStudy (anonymous):

this is way way (way) easier than the last one a) subtract 3 from both sides

OpenStudy (wilder.monday):

done :)

OpenStudy (anonymous):

rewrite in equivalent logarthmic form, i.e go from \[e^{3x+2}=20\] to \[3x+2=\ln(20)\] that is the only none algebra part

OpenStudy (anonymous):

"non algebra" part

OpenStudy (anonymous):

then solve as normal, subtract 2, then divide by 3

OpenStudy (wilder.monday):

thats it ?! easy!

OpenStudy (anonymous):

way way (way) easier than the last one you can just about do the last steps in your head you should have \[x=\frac{\ln(20)-2}{3}\]

OpenStudy (wilder.monday):

yep and i got 0.3319...

OpenStudy (anonymous):

if you are supposed to answer as a decimal, then that is probably correct

OpenStudy (wilder.monday):

yes decimals are fine

OpenStudy (wilder.monday):

can you help with two more?

OpenStudy (anonymous):

sure btw your answer is right

OpenStudy (wilder.monday):

okay thanks :)

OpenStudy (wilder.monday):

these next ones look long and hard

OpenStudy (wilder.monday):

and i just realized that sounded very wrong whoops

OpenStudy (anonymous):

oh great (maybe they are not so long and hard lol)

OpenStudy (wilder.monday):

maybe im over reacting lol

OpenStudy (wilder.monday):

log (2-1) = log (4x-3) - log x

OpenStudy (anonymous):

there must be a typo there right? is it really \[\ln(2-1)\] on the left?

OpenStudy (wilder.monday):

it says log(2-1)

OpenStudy (wilder.monday):

and then the rest on the other side

OpenStudy (wilder.monday):

is that bad

OpenStudy (anonymous):

ok if that is what it really says (seems weird) then it is pretty clear that \(2-1=1\)right?

OpenStudy (wilder.monday):

yes that seems about right lol

OpenStudy (anonymous):

so you really have \[\log (1) = \log (4x-3) - \log (x)\]